# Nicer Than Humans: How do Large Language Models Behave in the Prisoner’s Dilemma?

### v1

@@ 5 Discussion and Conclusion
Our study contributes to the broad literature on behavioral studies of Large Language Models as artificial social agents. Specifically, we study the responses of Llama2 to the prototypical social scenario of the Prisoner’s Dilemma. Building upon prior studies that explored the application of LLMs in classic Game Theory experiments, our work introduces a more systematic experimental setup that incorporates quantitative checks to better align the LLM responses to the complex task description. We have shown that aspects like prompt comprehension, memory representation, and duration of the simulation play crucial roles, with the potential to significantly distort the experimental outcomes, if left unchecked. Our framework provides a quantitative guide for selecting the simulation variables and improving the prompt.
Our findings add a new benchmark to the body of work that explored the outcomes of iterated games, both among humans and among AI agents. In contrast to the behavioral patterns observed in humans playing the Prisoner’s Dilemma game (Dal Bó and Fréchette 2011), Llama2 displayed a more marked propensity towards cooperation. Under conditions that disincentivize cooperation, most human players adopt a stance of complete defection, whereas the LLM’s strategy, albeit mostly uncooperative, is characterized by an initial trust in the opponent’s cooperation ($\mathsf{Nice}$), reminiscent of the strategy known in Game Theory as $\mathsf{Grim\ Trigger}$. When the environment is more favorable to cooperative play, human strategies often resemble either $\mathsf{Grim\ Trigger}$ or $\mathsf{Tit\ For\ Tat}$, while Llama2 tends towards a consistently cooperative approach. Notably, Llama2’s shift from $\mathsf{Grim\ Trigger}$ to $\mathsf{Always\ Cooperate}$ occurs quite abruptly when the probability of the opponent’s defection drops below 30%.
@@
However, it is important to acknowledge the limitations of our work, which open the way for further research. First, our analysis was conducted using a single model, specifically Llama2, which, at the time of writing, is one of the top-performing open-source models available (Touvron et al. 2023). Nevertheless, the field is progressing at an unprecedented pace, with new models being introduced regularly. To determine whether the behavioral patterns observed in our study are consistent across different models, it is important to conduct comparative analyses with models that vary significantly in terms of parameter size and the volume of their training data. Second, our study’s scope was limited to assessing the LLM’s responses to random strategies, and with a fixed payoff structure. Exploring the LLM’s interactions with more sophisticated opponents would enable us to better delineate the boundaries of LLMs’ inferential abilities in social contexts, and to draw more detailed behavioral profiles under a broader spectrum of conditions that more comprehensively represent prototypical social scenarios. Last, the experimental framework of our study considers only a single LLM agent. Creating social groups of AI agents that interact through iterated games like the Prisoner’s Dilemma would open up a wealth of opportunities to study emergent behaviors in synthetic societies, an avenue of research that is increasingly recognized as fundamental for a proper understanding of how LLMs can affect human societies (Bail 2024). Last, despite the numerical guidelines we implemented to evaluate the quality of the prompt, our refinements of the prompt were not guided by any principle other than experience and instinct. Our attempt to use Chain-of-Thought (Kojima et al. 2022; Zhou et al. 2023) as a structured way to approach prompt revision resulted in prompts that did not improve performance in our prompt comprehension question inventory. In this respect, our work provides yet another example of how prompt engineering would benefit from supporting tools to constrain its highly discretionary nature.
@@ 6.1 LLMs as Agents
Argyle et al. (2023) instructed LLMs to answer surveys as if they belonged to specified socio-demographic groups. They showed a high similarity between the responses generated by the LLM and those provided by the demographic groups it was asked to emulate. LLMs impersonating human agents with different profiles were used to create synthetic social networks, with the objective of observing emergent social behavior, most notably opinion dynamics and information spreading (Chuang et al. 2023; De Marzo, Pietronero, and Garcia 2023; He, Wallis, and Rathje 2023). Park et al. (2023) developed a society with synthetic agents interacting with elements of a synthetic world, showed that those agents were able to adopt behaviors that are typical of humans without being directly prompted to do so. Using the same framework, Ren et al. (2024) showed that those agents were also able to build and spread social norms.
@@ 6.2 LLMs in Game Theory
Early work on the application of LLMs to Game Theory experiments touched upon both 1-time and iterated games. Single-iteration experiments offer limited insight into LLM behavior. Brookins and DeBacker (2023) showed that LLMs are more biased towards fairness and cooperation when compared to a human baseline sample. In contrast, Aher, Arriaga, and Kalai (2023) found that there is an overall alignment between the LLM-based agent behavior and the human ones. When focusing on iterated games, the spectrum of patterns that can be identified expands, allowing more refined analysis. For example, Akata et al. (2023) managed to identify that LLMs can be particularly unforgiving. Mei et al. (2024) discovered instead that the same models show a higher cooperation rate than compared to humans. Fan et al. (2023) exploited the iteration of games to check the level of the opponent’s strategy that the LLM was able to infer from the history of actions. They showed that the inference capability of the LLM is limited, calling for more systematic approaches to structure memory and prompts.
7 Ethical Considerations
@@ Acknowledgments
NF acknowledges the support from the Danish Data Science Academy through the DDSA Visit Grant (Grant ID: 2023-1856) and from Politecnico di Milano through the scholarship “Tesi all’estero a.a. 2023/24-Primo bando”. FP acknowledges support from the European Union (NextGenerationEU project PNRR-PE-AI FAIR) and the Italian Ministry of Education (PRIN PNRR grant CODE prot. P2022AKRZ9 and PRIN grant DEMON prot. 2022BAXSPY). LMA acknowledges the support from the Carlsberg Foundation through the COCOONS project (CF21-0432).
@@
Figures 6, 7 and 8 show examples of the three parts of the final prompt we used in our experiments.
@@
Figure 6: System prompt describing the players’ roles, the rules of the games, the number of rounds to be played, and the player’s objective. This part of the prompt is fixed in all simulations.
@@
Figure 7: Contextual prompt containing information about: the game history in the last $n$ rounds ($5$ in this example), the overall amount of times each player chose each action, the overall amount of points collected by each player, and the current round. For each turn, the prompt contains: the action played by each player and the points collected by each player. This prompt changes at each round.
@@
Figure 8: Instructing prompt. The LLM is instructed on the nature and format of the answer. This part of the prompt component is replaced with prompt comprehension questions in the phase of prompt tuning.

### v2

@@ 5 Discussion and Conclusion
Our study contributes to the broad literature on behavioral studies of Large Language Models as artificial social agents. Specifically, we study the responses of Llama2, Llama3, and GPT3.5 to the prototypical social scenario of the Prisoner’s Dilemma. Building upon prior studies that explored the application of LLMs in classic Game Theory experiments, our work introduces a more systematic experimental setup that incorporates quantitative checks to better align the LLMs’ responses to the complex task description. We have shown that aspects like prompt comprehension, memory representation, and duration of the simulation play crucial roles, especially for less powerful models, with the potential to significantly distort the experimental outcomes, if left unchecked. Our framework provides a quantitative guide for selecting the simulation variables and improving the prompt.
Our findings add a new benchmark to the body of work that explored the outcomes of iterated games, both among humans and among AI agents. In contrast to the behavioral patterns observed in humans playing the Prisoner’s Dilemma game (Dal Bó and Fréchette 2011), the three models displayed a more marked overall propensity towards cooperation. Under conditions that disincentivize cooperation, most human players adopt a stance of complete defection, whereas the LLMs’ strategy, albeit mostly uncooperative, is characterized by an initial trust in the opponent’s cooperation ($\mathsf{Nice}$), reminiscent of the strategy known in Game Theory as $\mathsf{Grim\ Trigger}$. When the environment is more favorable to cooperative play, Llama3 playing $\mathsf{Grim\ Trigger}$ becomes comparable to human strategies that often resemble either $\mathsf{Grim\ Trigger}$ or $\mathsf{Tit\ For\ Tat}$. Differently, Llama2 and GPT3.5 tend towards a consistently cooperative approach. Notably, Llama2’s shift from $\mathsf{Grim\ Trigger}$ to $\mathsf{Always\ Cooperate}$ occurs quite abruptly as the opponent’s defection probability drops below 30%, while GPT3.5’s transition is smoother and starts when $\alpha$ exceeds $0.2$.
@@
Overall, our findings offer a robust baseline for understanding LLMs’ behaviors in the Iterated Prisoner’s Dilemma (IPD), a widely used Game Theory experiment for assessing agents’ responses to conflictual scenarios. Establishing this baseline allows for clearer differentiation between the inherent tendencies of the models and the effects of specific elements within the game setting.
However, it is important to acknowledge the limitations of our work, which open the way for further research. First, our analysis was conducted using three models, specifically Llama2, Llama3, and GPT3.5, which, at the time of writing, are among the most advanced models available (Touvron et al. 2023; Brown et al. 2020) [^5] https://ai.meta.com/blog/meta-llama-3/. Nevertheless, the field is progressing at an unprecedented pace, with new models being introduced regularly. To determine whether the behavioral patterns observed in our study are consistent across many different models, it is important to conduct comparative analyses with models that vary significantly in terms of parameter size and the volume of their training data. Second, our study’s scope was limited to assessing the LLM’s responses to random strategies, and with a fixed payoff structure. Exploring the LLM’s interactions with more sophisticated opponents would enable us to better delineate the boundaries of LLMs’ inferential abilities in social contexts, and to draw more detailed behavioral profiles under a broader spectrum of conditions that more comprehensively represent prototypical social scenarios. We explored only the use of Zero-Shot Chain-of-Thought techniques without obtaining any improvement. Other options like Tree-of-Thought (Yao et al. 2023) techniques or external modules (Wang et al. 2024) can be employed to test different reasoning conditions for the models. Furthermore, the experimental framework of our study considers only a single LLM agent. Creating social groups of AI agents that interact through iterated games like the Prisoner’s Dilemma would open up a wealth of opportunities to study emergent behaviors in synthetic societies, an avenue of research that is increasingly recognized as fundamental for a proper understanding of how LLMs can affect human societies (Bail 2024). Last, despite the numerical guidelines we implemented to evaluate the quality of the prompt, our refinements of the prompt were not guided by any principle other than experience and instinct. Our attempt to use Chain-of-Thought (Kojima et al. 2022; Zhou et al. 2023) as a structured way to approach prompt revision resulted in prompts that did not improve performance in our prompt comprehension question inventory (see Figure A6 in Appendix). In this respect, our work provides yet another example of how prompt engineering would benefit from supporting tools to constrain its highly discretionary nature. Another interesting direction for future research is to explore persona prompting (Hu and Collier 2024), where models are guided towards specific behaviors, such as altruism or selfishness. However, in this study, we intentionally avoided using personas to influence the outputs of the LLMs, as our primary objective is to assess the models’ inherent biases towards different behavioral patterns in game-theoretical scenarios. Similarly, models could be enhanced by incorporating a planning module that considers the potential outcomes of various actions over multiple steps (Kambhampati et al. 2024).
@@ 6.1 LLMs as Agents
Argyle et al. (2023) instructed LLMs to answer surveys as if they belonged to specified socio-demographic groups. They showed a high similarity between the responses generated by the LLM and those provided by the demographic groups it was asked to emulate. LLMs impersonating human agents with different profiles were used to explore the negotiation abilities of the models (Davidson et al. 2024) or to create synthetic social networks, to observe emergent social behavior, most notably opinion dynamics and information spreading (Chuang et al. 2023; De Marzo, Pietronero, and Garcia 2023; He, Wallis, and Rathje 2023). Park et al. (2023) developed a society with synthetic agents interacting with elements of a synthetic world, showed that those agents were able to adopt behaviors that are typical of humans without being directly prompted to do so. Using the same framework, Ren et al. (2024) showed that those agents were also able to build and spread social norms, while Piatti et al. (2024) developed a similar framework to investigate the robustness of LLMs societies.
@@ 6.2 LLMs in Game Theory
Early work on the application of LLMs to Game Theory experiments touched upon both 1-time and iterated games. Single-iteration experiments offer limited insight into LLM behavior. Brookins and DeBacker (2023) showed that LLMs are more biased towards fairness and cooperation when compared to a human baseline sample. In contrast, Aher, Arriaga, and Kalai (2023) found that there is an overall alignment between the LLM-based agent behavior and the human ones. Investigating the ability of LLMs to predict human choices in 1-time games, Capraro, Di Paolo, and Pizziol (2024) showed that only more powerful models are capable of doing it, although they overestimate the altruistic tendency of human players. When focusing on iterated games, the spectrum of patterns that can be identified expands, allowing more refined analysis. For example, Akata et al. (2023) managed to identify that LLMs can be particularly unforgiving. Mei et al. (2024) discovered instead that the same models show a higher cooperation rate than compared to humans. Fan et al. (2023) exploited the iteration of games to check the level of the opponent’s strategy that the LLM was able to infer from the history of actions. They showed that the inference capability of the LLM is limited, calling for more systematic approaches to structure memory and prompts. Testing a novel benchmark with different games, Duan et al. (2024) studied multiple LLMs showing that closed-source models tend to achieve better performance than open-source ones. Although they did not examine the behavioral characteristics of the LLMs’ responses, they also found that Zero-Shot Chain-of-Thought techniques are not always beneficial, which aligns with our results.
Ethical Considerations
@@
Code and Data Availability
All code, prompts, and game traces will be made available on GitHub.
@@ Acknowledgments
This work was partially supported by the Italian Ministry of Education ( PRIN grant DEMON prot. 2022BAXSPY) and the European Union (NextGenerationEU project PNRR-PE-AI FAIR). NF acknowledges the support from the Danish Data Science Academy through the DDSA Visit Grant (Grant ID: 2023-1856) and from Politecnico di Milano through the scholarship “Tesi all’estero a.a. 2023/24-Primo bando”. LMA acknowledges the support from the Carlsberg Foundation through the COCOONS project (CF21-0432).
@@
Prompts and their Variations
Figures A1, A2 and A3 show examples of the three parts of the final prompt we used in our experiments. Figures A4 and A5 respectively show the variation in the instructing prompt when using the Zero-Shot Chain-of-Thought techniques from Kojima et al. (2022) and Zhou et al. (2023). Figure A6 displays the accuracy obtain by Llama2 in the comprehension questions when using the two variations and the unvaried prompt.
Figure A7 shows the comparison in the accuracy of Llama2 between the initial prompt without explicit score counts and the final version of it.
@@
Figure A1: System prompt describing the players’ roles, the rules of the games, the number of rounds to be played, and the player’s objective. This part of the prompt is fixed in all simulations.
@@
Figure A2: Contextual prompt containing information about: the game history in the last $n$ rounds ($5$ in this example), the overall amount of times each player chose each action, the overall amount of points collected by each player, and the current round. For each turn, the prompt contains: the action played by each player and the points collected by each player. This prompt changes at each round.
@@
Figure A3: Instructing prompt. The LLM is instructed on the nature and format of the answer. This part of the prompt component is replaced with prompt comprehension questions in the phase of prompt tuning.
Instructing prompt
Remember to use only the following JSON format:
{"action": <ACTION_of_A>, "reason": <YOUR_REASON>} Answer saying which action player A should play. Remember to answer using the right format.[/INST]
Let’s think step by step
Figure A4: Variation of the instructing prompt with Kojima et al. (2022) Zero-Shot Chain-of-Thought.
Instructing prompt
Remember to use only the following JSON format:
{"action": <ACTION_of_A>, "reason": <YOUR_REASON>} Answer saying which action player A should play. Remember to answer using the right format.[/INST]
Let’s work this out in a step-by-step way to be sure we have the right answer
Figure A5: Variation of the instructing prompt with Zhou et al. (2023) Zero-Shot Chain-of-Thought.
Figure A6: Accuracy of Llama2’s responses to the prompt comprehension questions defined in Table 1 using two Zero-Shot Chain-of-Thought variations of the prompt compared to the accuracy obtained with the unvaried prompt. We show 95% confidence intervals, computed from 100 games.
Figure A7: Accuracy of Llama2 responses to the prompt comprehension questions defined in Table 1. The accuracy of the final full prompt is compared with an earlier version of the prompt that lacked a summary of the cumulative rewards achieved by the players. We show 95% confidence intervals, computed from 100 games.
Effect of Memory Window Size
Figure A8 shows the probability of cooperation for Llama3 and GPT3.5 when using memory windows of different sizes. Different from Llama2, neither of these models significantly changes its behavior depending on the window size. In particular, the trend observed for GPT3.5, which tends to stabilize around the value 0.2, is in accordance with the average $p_{coop}$ shown in Figure 5 for $\alpha=0$
Figure A8: Llama3’s and GPT3.5’s probability of cooperation ($p_{coop}$) against an Always Defect opponent, when using a memory window size of 10 vs. including the full game history in the prompt.
Effect of Temperature
We explore the impact of varying the temperature hyperparameter on the probability of cooperation $p_{coop}$ of Llama3 and GPT3.5111HuggingFace discontinued access to Llama2 through their Inference API in May 2024. Since then, the model has been available only dedicated server hosting at considerable costs. We therefore limit our tests on temperature to Llama3 and GPT3.5 only.. For each model and each temperature ($0.1$ and $1.0$), we run 10 games of 100 rounds each against opponents with varying cooperation probability $\alpha$. Figure A9 presents the average probability of cooperation for Llama3 and GPT3.5 for three different temperature values up to $1.0$. At temperatures greater than $1.0$, the models tend to produce seemingly random tokens, which renders their output unusable.
The Pearson correlation between the values of $p_{coop}$ across different temperature pairs is in the range $[0.97-1]$, indicating that the temperature does not affect the general trends of $p_{coop}$ as $\alpha$ varies. However, the growth of the cooperation curve for GPT3.5 turns from roughly linear to a sigmoid as temperature decreases. This suggests that, in some models, different levels of determinism can influence the boundaries of their decision states. In this specific case, it appears that the noise that the increased temperature adds to the generated choices smoothens the sharp transition between the uncooperative and cooperative states that can be observed at low temperatures.
Figure A9: Llama3’s and GPT3.5’s probability of cooperation ($p_{coop}$) against $\mathsf{Unfair\ Random}$ opponents with increasing cooperation probability $\alpha$ for different values of the temperature hyperparameter. We show 95% confidence intervals, computed from 10 games.